OpenStudy (deercult):

Solve ln 5 + ln (2x) = 5. Round your answer to the nearest hundredth.

OpenStudy (loviy):

First you need to isolate the term with an x, so the idea here is to get rid of the ln5 on left side. Take note that ln5 is being added on left side, so we need to use the opposite operation of Addition, which is?

OpenStudy (deercult):

Subtraction. :>

OpenStudy (sweetburger):

You can also combine the natural logs via \[\ln(a)+\ln(b)=\ln(a \times b)\]

OpenStudy (loviy):

Yes, wait.. we could use the property: lna + lnb = ln(ab) first.

OpenStudy (loviy):

For example ln3 + ln(4x) = ln12x..follow?

OpenStudy (deercult):

I follow! So it'd be ln 5 + ln(2x) = ln10x?

OpenStudy (sweetburger):

yes correct

OpenStudy (loviy):

Yes, exactly!

OpenStudy (loviy):

Then, we could use the opposite of natural logarithm..

OpenStudy (loviy):

Take note that e^(lnu) = u.

OpenStudy (loviy):

So far our equation is ln(10x) = 5. For example if we have ln(7x) = 4: e^(ln(7x)) = e^4 ---> 7x = e^4.

OpenStudy (loviy):

Does that make sense?

OpenStudy (deercult):

I think so, yes. 10x = e^5, then?

OpenStudy (loviy):

Exactly!

OpenStudy (loviy):

Then, we just need to get rid of the 10 on left side, what is the opposite operation of multiplication?

OpenStudy (deercult):

Division!

OpenStudy (loviy):

Yes, so what will we have if we divide 10 on both sides?

OpenStudy (deercult):

x = (e^5)/10

OpenStudy (loviy):

Yes, and if we will calculate that and round to nearest hundredth?

OpenStudy (deercult):

Which is 14.84!

OpenStudy (loviy):

Perfect! Good job!

OpenStudy (deercult):

Thanks! You were a lot of help!

OpenStudy (loviy):

Pleasure to help you! God bless you!

OpenStudy (deercult):

You, too! Have a great night!

OpenStudy (loviy):

You too!